2/5w - w = -1/5(3w + 2)
I got 0w=-2
that can't be right?

Answer by Alan3354(69443)

If you mean:
(2/5)w - w = (-1/5)(3w + 2)
Multiply by 5
2w - 5w = -(3w + 2) = -3w - 2
-3w = -3w - 2
0 = -2
----------------------
--> inconsistent, no solution
There is no value of w that will make the equation true.